SOHO Intercalibration Joint Observing Programme 03

INTERCALIBRATION 3

Author(s): R.A. Harrison (CDS), J.-P. Delaboudiniere (EIT), D. Judge (SEM) Progress:

Draft Scheme  February 2 1995
Distributed to Relevant Parties  February 2 1995

Objective: Intensity cross-calibration using full Sun images. Conditions Necessary to Run: Requires relatively quiet Sun - to minimise changes during CDS full Sun raster. Must include CDS and SEM at a minimum. Method: The idea of this scheme is that CDS performs a full Sun raster at 304Å for comparison to the SEM integrated Sun 304Å monitor. EIT should be used both for cross calibration and to check for any transient features. This requires EIT enhanced telemetry before and after the CDS scans. Total duration about 45 minutes. Pointing: This scheme makes use of the CDS AERON Study which uses the NIS and rasters the entire instrument over the full disc. Frequency: This scheme should be run periodically, but rarely, since it involves 130 repointings of CDS! Operating Details:

• CDS

GIS, with 90x240 arcsec slit. 5 second exposure at each location. CDS ”rastered” over full disc - 130 repointings. Total duration 13 minutes. Extract GIS band 2.

CDS Study ID: AERON

• EIT

Produce full Sun image in 304Å using enhanced EIT telemetry mode just prior to and after the CDS scans. 15 minutes in each case assuming 10:1 compression.

• SEM

Monitor full Sun 304Å signal through out the 45 minute period.